sunisround:

Given a degree 5 polynomial with real coefficients has roots 3,2-i,6+i, find the other missing roots?

sunisround:

I got 2+i and 6-i as the other missing roots am i correct?

Vocaloid:

that's correct, good work

sunisround:

ok because i recalled that law which you said that a-bi can be a+bi something like that correct me if i am wrong.

Vocaloid:

yes, that's the conjugate root theorem

sunisround:

alright thank you.

sillybilly123:

awesome teaching !!

sillybilly123:

easy to say with hindsight: \((x - z_0 ) ( x - \bar z_0) = x^2 - x (\bar z_0 + z_0) + z_0 \bar z_0 = x^2 - 2 x_0 x + |z_0|^2\) All of which follows from the fact that \(z_0 + \bar z_0 = 2 x_0\) and \(z_0 \bar z_0 = |z_0|^2\) So: \((x - \alpha)(x - \beta)(x - z_0 ) ( x - \bar z_0) = (x - \alpha)(x - \beta) ( x^2 - 2 x_0 x + |z_0|^2)\)